A gardener is fencing off a rectangular area with a fixed perimeter of 88 ft. What is the maximum​ area?
Answer #1

BB Perimeter 2 (4+ B ) = BB Lt B = 44 B = 44-L LXB Area A L (44-L ALL) = 44L-12 To maximize Area set Area set A(L) = 0 ACL)